OpenStudy (anonymous):

A collection of dimes is arranged in a triangular array with 18 coins in the base row, 17 in the next, 16 in the next, and so forth with 1 dime in the last row. Find the value of the collection.

OpenStudy (dumbcow):

first find num of dimes then multiply by 0.1 to get value in dollars sequence goes 18,17,16...1 so you are subtracting 1 each time there is a formula for sum of arithmetic sequence sum = 1/2*n*(a1+an) n=num of rows

OpenStudy (anonymous):

im still a little lost

OpenStudy (dumbcow):

ok just concentrate on the formula for the sum of the series we need a value for n, a1, and an n is num of rows a1 is first num in sequence an is last num in sequence

OpenStudy (anonymous):

hmmmm

OpenStudy (dumbcow):

whats first num in sequence

OpenStudy (dumbcow):

how many dimes in first row? how many dimes in last row?

OpenStudy (dumbcow):

still lost? ok top row has one dime bottom row has 18 dimes since each row has one more than prev row there are a total of 18 rows 1,2,3,4,5,6,7,8,9,10,11,12,13,14,15,16,17,18 sum = 1/2*18*(1+18) sum = 1/2*18*19 sum = 9*19 sum = 171 171 dimes = 17.10 in dollars

OpenStudy (anonymous):

ah ok got it... thank you!!!

OpenStudy (dumbcow):

no prob